40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 169 of 336  Not logged in ELOG logo
ID Date Author Type Category Subject
  8445   Thu Apr 11 16:29:23 2013 AlbertUpdateoptical tablesOptical Table Toolboxes Update

Quote:

 There are some tips for how to appy nail polish on YouTube from MKNails and MissJenFABULOUS. Their tips on how to prepare the site for a strong bonding strength are probably helpful for our gold/nickel coated tools. For chrome tools we may need to abrade the surface with a stone or fine sandpaper for it to take the layer better. IF the YouTube videos don't do it for you, then I suggest contacting Tom Evans at LLO to find out what kind of nail polish he uses.

 This is the tentative box placement per optical table. The toolboxes are going to be color-coded by a combination of two colors (the order won't matter). The side of each toolbox will have a little panel to let you know which box corresponds to which set of colors.

On the diagram, the set of colors is simply the color of the box border and the color of the text.

If anyone has a problem with any of the colors or the box placement let me know before they are installed and become an annoyance:

Diagram.png

Box Placements:

ETMY: Box will be attached to the underside of the table by magnets. The box will be on the north side of the optical table.

POY: Box will be attached to the side of the optical table by magnets. The box will be on the west side of the optical table.

BSPRM: Box will be attached to the side of the optical table by magnets. The box will be on the west side of the optical table.

AS: Box will be attached to the side of the optical table by magnets. The box will be on the north side of the optical table.

PSL1: Box will be inside the optical table, in the northeast corner.

PSL2: Box will be inside the optical table, in the southwest corner.

POX: Box will be attached to the side of the optical table by magnets. The box will be on the south side of the optical table.

MC2: Box will be attached to the side of the optical table by magnets. The box will be on the south side of the optical table.

ETMX: Box will be attached to the side of the optical table by magnets. The box will be on the east side of the optical table.

  8444   Thu Apr 11 11:58:21 2013 JenneUpdateComputersLSC whitening c-code ready

The big hold-up with getting the LSC whitening triggering ready has been a problem with running the c-code on the front end models.  That problem has now been solved (Thanks Alex!), so I can move forward.

The background:

We want the RFPD whitening filters to be OFF while in acquisition mode, but after we lock, we want to turn the analog whitening (and the digital compensation) ON.  The difference between this and the other DoF and filter module triggers is that we must parse the input matrix to see which PD is being used for locking at that time.  It is the c-code that parses this matrix that has been causing trouble.  I have been testing this code on the c1tst.mdl, which runs on the Y-end computer.  Every time I tried to compile and run the c1tst model, the entire Y-end computer would crash.

The solution:

Alex came over to look at things with Jamie and me.  In the 2.5 version of the RCG (which we are still using), there is an optimization flag "-O3" in the make file.  This optimization, while it can make models run a little faster, has been known in the past to cause problems.  Here at the 40m, our make files had an if-statement, so that the c1pem model would compile using the "-O" optimization flag instead, so clearly we had seen the problem here before, probably when Masha was here and running the neural network code on the pem model.  In the RCG 2.6 release, all models are compiled using the "-O" flag.  We tried compiling the c1tst model with this "-O" optimization, and the model started and the computer is just fine.  This solved the problem.

Since we are going to upgrade to RCG 2.6 in the near-ish future anyway, Alex changed our make files so that all models will now compile with the "-O" flagWe should monitor other models when we recompile them, to make sure none of them start running long with the different optimization. 

The future:

Implement LSC whitening triggering!

  8443   Thu Apr 11 10:15:55 2013 SteveUpdateLockingPRM yaw oplev transferfunction

See   Feb 2012 PRM yaw transferfunctions, also check Valera's modified  side sensor may effect yaw motion

 

  8442   Thu Apr 11 03:38:40 2013 DenUpdateLockingangular motion

Spectra of BS, PRM, ITMX, ITMY are attached with oplevs ON and OFF (in units of urad). Loops reduce RMS from ~2urad to ~0.3urad but phase margin should be increased. REF traces show loop OFF. <-- really?

Note how PRM pitch and yaw spectra are different in the frequency range 0.5 - 7 Hz; yaw is factor of 50 larger then pitch at 2 Hz.

Attachment 1: oplevs.pdf
oplevs.pdf oplevs.pdf oplevs.pdf oplevs.pdf
  8441   Thu Apr 11 03:25:29 2013 JenneHowToSUSIdea for how to properly balance SUS actuators
We have calibrated the overall actuators of each suspension independent of the optical levers. So, we know how much we are 
moving the optic in POS in real units as a result of the dither we inject for the lockin measurement. The amount the oplev beam 
appears to move if there is only POS motion is
d/cos(theta)
where theta is the oplev's angle of incidence and d is the distance the optic has moved in POS.  None of the of the steering mirrors in the 
oplev path matter. 

I propose that I will add an option in the lockin path to subtract away the apparent angle from the oplev output just before the signal 
goes into the lockin module.  Then we will be balancing the actuators based on only the actual angular motion.

The success of this technique depends on how well we know our actuator calibration and the oplev angle of incidence. This also 
assumes that the oplev beam is centered on the optic, so we don't have beam displacement from A2L of the oplev beam, which then 
makes another apparent angular motion.  I suspect that we are close enough that we won't have to worry about this effect.
  8440   Thu Apr 11 03:23:12 2013 DenUpdateGeneralMCL threshold

MC down script is too slow to block MC_L when the cavity goes out of lock. As a result the loop strongly kicks MC2. We decided to make a threshold inside MCS model on MC TRANS that will block MC_L during lock loss. This is a lower threshold. Upper threshold can be slow and is implemented inside MC up script.

Fast threshold can be set inside MC2 POS. I did not correct MC2 top level medm screen as it is the same for all core optics.

Note: Fast trigger will also block ALS signal if MC loose lock.

  8439   Thu Apr 11 02:49:18 2013 DenUpdateLockingPRCL on carrier

Jenne, Den

We suspect PRM shows significant length to angle coupling due to large oplev beam angle in yaw.  Tonight we locked PRCL with ITMs.

We could lock PRCL on carrier to power recycling gain of 15. Lock continued for a few hours but power rin RMS was 0.15.

We triggered and normalized on POP_DC. MICH gain was -1 (filters FM3-5), PRCL gain was -8 (filters FM2,4,5,6,9).

MC_L was OFF during locking.

 

Attachment 1: pop_rin.pdf
pop_rin.pdf
Attachment 2: power.png
power.png
  8438   Thu Apr 11 02:00:21 2013 JenneUpdateLockingTRY gone???

TRY signals are all gone!  Both the PD and the camera show no signal.  I went down there to turn off the lights, and look to see what was up, and I don't see any obvious things blocking the beam path on the table.  However, Steve has experimentally bungeed the lids down, so I didn't open the box to really look to see what the story is.

Absent TRY, I redid the IFO alignment.  Yarm locked, so I assumed it was close enough.  I redid Xarm alignment pretty significantly.  Transmission was ~0.5, which I got up to ~0.85 (which isn't too bad, since the PMC transmission is 0.74 instead of the usual 0.83).  I then aligned MICH, and PRM.  After fixing up the BS alignment, the POP beam wasn't hitting the POP PD in the center any more.  I centered the beam on the PD, although as Gabriele pointed out to me a week or two ago, we really need to put a lens in front of POP, since the beam is so big.  We're never getting the full beam when the cavity flashes, which is not so good.

Den is still working on locking, so I'll let him write the main locking report for the night.

We see that the PRC carrier lock seems to be more stable when we lock MICH with +1 for ITMY and -1 for ITMX, and PRCL with -1 for both ITMs.  This indicates that we need to revisit the systematic problem with using the PRM oplev to balance the coils, since that oplev has a relatively wide opening angle.  I am working on how to do this.

  8437   Wed Apr 10 15:49:22 2013 AnnalisaConfigurationCOMSOL TipsYend table eigenfrequency simulation with COMSOL

 I made a Simulation with COMSOL for the Yend table. Mainly, I tried to see how the lower eigenmode changes with the number and the size of the posts inside.

The lateral frame is just sitting on the table, it is fixed by its weight. I also put a couple of screws to fix it better, but the resulting eigenfrequency didn't change so much (less than 1 Hz). 

In Fig. 1 I didn't put any post. Of course, the lowest eigenfrequency is very low (around 80 Hz).

Then I added 2 posts, one per side (Fig. 2 and Fig. 3), with different diameter.

In some cases posts don't have a base, but they are fixed to the table only by a screw. It is just a condition to keep them fixed to the table

Eventually I put 4 posts, 2 per side. 

The lowest eigenfrequency is always increasing.

At the end I also put a simulation for 4 1.6 inch diameter posts without base, and the eigenfrequency is slightly higher. I want to check it again, because I would expect that the configuration shown in Fig.5a could be more stable.

P.S.: All the post are stainless steel.

 

Attachment 1: Pics_end_table.pdf
Pics_end_table.pdf Pics_end_table.pdf Pics_end_table.pdf
  8436   Wed Apr 10 14:25:04 2013 ManasaUpdate40m UpgradingEndtable upgrade for auxiliary green laser : LWE Alberto laser beam profile

I measured the beam profile of the Alberto laser (that will be the ETMY end laser) as I found the data inconsistent with the measurements done earlier.

Method

The laser was set to nominal current (ADJ =0) and the output measured 330mW. I used a 99% BS and measured the beam profile of the transmitted light at several points along propagation using Beamscan. I am attaching the data and matlab script for the fit for future reference.

LWE_Alberto_BP.png

Attachment 2: beam_waist.zip
  8435   Wed Apr 10 07:37:33 2013 SteveUpdatePSLPMC locked

Quote:

Quote:

The PMC locked manually. MC  grabbed lock instantaniously

 PMC locked

 PMC locked.

Attachment 1: pmc4d.png
pmc4d.png
  8434   Wed Apr 10 03:59:41 2013 DenConfigurationIOOTurn on MCL

Quote:

 My belief is that the frequency noise from the unstabilized MC is making the PRC locking harder. This will be investigated by tuning the shape of the MCL/MCF crossover so that we can turn it on without ruining the arm cavity spectra. Since the PRC length is ~2x smaller than the MC, we would expect it to be less sensitive to the MC frequency noise. But, since there is some common mode rejection in there, this may not be true. We'll only know by measuring PRC control signal with MCL on/off.

 I think if we make MCL UGF higher then 20 Hz, arm cavity spectra will feel it. It might be possible to use a combination of feedback and feedforward control from ground seismometers. I made MCL UGF at 3 Hz to reduce 1 Hz motion of the pendulum; feedforward OAF subtracted the stack at 3.3 Hz. Once OAF converged, I blocked adaptation and the filter became static FIR. MC length RMS was reduced by a factor of 10 and arm cavity spectra was not affected at frequencies >20 and became better at low frequencies. We'll see if this enough.

On the attached plot red color shows MC_F with MC_L OFF, blue - MC_L is ON, green - MC_L and OAF are ON.

Then I locked PRCL (using AS_Q and REFL55_I) to carrier and aligned the cavity. Power RIN was 50-70% and 00 beam on the POP camera was moving significantly. BS oplev was shaking the optics at 5 Hz. I fixed it, but there should be something else as RIN was still high.

Attachment 1: MCL.pdf
MCL.pdf
  8433   Wed Apr 10 01:10:22 2013 JenneUpdateLockingConfigure screen and scripts updated

I have gone through the different individual degrees of freedom on the IFO_CONFIGURE screen (I haven't done anything to the full IFO yet), and updated the burt snapshot request files to include all of the trigger thresholds (the DoF triggers were there, but the FM triggers and the FM mask - which filter modules to trigger - were not).  I also made all of the restore scripts (which does the burt restore for all those settings) the same.  They were widely different, rather than just different optics chosen for misaligning and restoring.

Before doing any of this work, I moved the whole folder ..../caltech/c1/burt/c1ifoconfigure to ..../caltech/c1/burt/c1ifoconfigure_OLD_but_SAVE , so we can go back and look at the past settings, if we need to.

I also changed the "C1save{DoF}" scripts to ask for keyboard input, and then added them as options to the CONFIGURE screen.  The keyboard input is so that people randomly pushing the buttons don't overwrite our saved burt files.  Here's the secret:  It asks if you are REALLY sure you want to save the configuration.  If you are, type the word "really", then hit enter (as in yes, I am really sure).  Any other answer, and the script will tell you that it is quitting without saving.

I have also removed the "PRM" option, since we really only want the "PRMsb" for almost all purposes.

Also, I removed access to the very, very old text file about how to lock from the screen.  That information is now on the wiki:  https://wiki-40m.ligo.caltech.edu/How_To/Lock_the_Interferometer

I have noted in the drop-down menus that the "align" functions are not yet working.  I know that Den has gotten at least one of the arms' ASSes working today, so once those scripts are ready, we can call them from the configure screen.

Anyhow, the IFO_CONFIGURE screen should be back to being useful!

  8432   Tue Apr 9 21:27:48 2013 JenneUpdate40m Upgrading TRY temporarily in place

Quote:

I've used a Y1 mirror to steer the Y transmission to an R98% BS. The reflected beam falls on PDA520 and the transmitted beam is steered to the camera. The earlier normalization of TRY is no more valid as the power distribution at the PD has changed.

 To take this into account, last night, I reduced the TRY gain by a factor of 2.  This is not exactly correct - when the layout is finalized we need to figure out what the pickoff situation used to be (we think, based on the Xend, that it could have been 0.5*0.9), and do the correct normalization.

  8431   Tue Apr 9 14:55:13 2013 JamieUpdateCDSoverbooked test points cause of DAQ problems

Folks were complaining that they were getting zeros whenever they tried to open fast channels in DTT or Dataviewer.  It turned out that the problem was that all available test points were in use in the c1lsc model:

lsc-gds.png

There is a limit to how many test points can be open to a single model (in point of fact I think the limit is on the data rate from the model to the frame builder, not the actual number of open test points).  In any event, they was all used up.  The grid at the bottom right of the C1LSC GDS screen was all full of non-zeros, and the FE TRATE number was red, indicating that the data rate from this model had surpassed threshold.

The result of this overbooking is that any new test points just get zeros.  This is a pretty dumb failure mode (ideally one would not be able to request the TP at all with an appropriate error message), but it is what it is.  This usually means that there are too many dtt/dataviewers left with open connections.

We tried killing all the open processes that we could find that might be holding open test points, but that didn't seem to clear them up.  Stuck open test points is another known problem.  Referencing the solution in #6968 I opened the diag shell and killed all test points everywhere:

controls@pianosa:~ 0$ diag -l -z
Set new test FFT
NDS version = 12
supported capabilities: testing  testpoints  awg  
diag> tp clear * *
test point cleared
diag> quit
EXIT KERNEL
controls@pianosa:~ 0$
  8430   Tue Apr 9 08:37:54 2013 SteveUpdate40m UpgradingEndtable upgrade for auxiliary green laser : TRY temporarily in place

Quote:

The TRY path on the end table is temporarily in place to help IFO locking.
The Y arm transmission was steered to get TRY back on the PD and the camera. I found that TRY is a couple of inches off in yaw at the end table (comparing to the CAD layout and the earlier layout) and I believe it is because of the changes in input pointing.
I've used a Y1 mirror to steer the Y transmission to an R98% BS. The reflected beam falls on PDA520 and the transmitted beam is steered to the camera. The earlier normalization of TRY is no more valid as the power distribution at the PD has changed.

Temporary  acrylic wind guard added between enclosure and ETMY transmission window to help IFO locking

  8429   Tue Apr 9 07:48:10 2013 SteveUpdatePSLPMC locked

Quote:

The PMC locked manually. MC  grabbed lock instantaniously

 PMC locked

Attachment 1: pmc10d.png
pmc10d.png
  8428   Tue Apr 9 01:46:40 2013 ZachUpdateGeneralRestarted elog

Again.

Quote:

with the script, as it was down.

 

  8427   Tue Apr 9 00:32:57 2013 ManasaUpdate40m UpgradingEndtable upgrade for auxiliary green laser : TRY temporarily in place

The TRY path on the end table is temporarily in place to help IFO locking.
The Y arm transmission was steered to get TRY back on the PD and the camera. I found that TRY is a couple of inches off in yaw at the end table (comparing to the CAD layout and the earlier layout) and I believe it is because of the changes in input pointing.
I've used a Y1 mirror to steer the Y transmission to an R98% BS. The reflected beam falls on PDA520 and the transmitted beam is steered to the camera. The earlier normalization of TRY is no more valid as the power distribution at the PD has changed.

  8426   Tue Apr 9 00:32:39 2013 ranaConfigurationIOOTurn on MCL

 Listening to the free swinging Y-arm, its clear that the fringe velocity is higher with the MCL path off, than on. I checked that the default gain of -300 was too high and changed it to -100 in the mcup script. With the higher gain value, there was clear gain peaking at just under 60 Hz (where the 60 Hz comb filter starts). We basically want the UGF to be between 20 and 60 Hz so that we can have the Bounce mode RG at 16.25 Hz and also the notch filter at 60 Hz.

Den is measuring and setting the UGF to be ~45 Hz.

With the MCL path on there is a high frequency horn-like noise in the Yarm when it locks. Its reduced a little by reducing the loop gain, but clearly this is just noise introduced by the MCL loop as Den noted before (and also tonight). He is cleaning up the whole MCL MEDM situation so that its useable by us. At the moment I've re-enabled it in the mcup.

My belief is that the frequency noise from the unstabilized MC is making the PRC locking harder. This will be investigated by tuning the shape of the MCL/MCF crossover so that we can turn it on without ruining the arm cavity spectra. Since the PRC length is ~2x smaller than the MC, we would expect it to be less sensitive to the MC frequency noise. But, since there is some common mode rejection in there, this may not be true. We'll only know by measuring PRC control signal with MCL on/off.

  8425   Tue Apr 9 00:15:18 2013 ManasaUpdate40m UpgradingEndtable upgrade for auxiliary green laser : populating the table

[Den, Annalisa, Manasa]

The Alberto laser was moved from the PSL table. The height of the heat sink rendered a beam height of only 3 inches. I did not want to deal with changing beam height at the table. So, we went ahead and used the old heat sink. I used the beam scan to make measurements of the beam width to match my mode-matching calculations and found some mismatch with the measurements done earlier. So I will measure the beam width again before alignment.

I will also have to change the layout because of the supporting posts that have come up with the new box. Annalisa is doing a COMSOL model to check what the thickness of these supporting posts should be so that the box stays stiff.

  8424   Mon Apr 8 22:43:44 2013 ranaUpdatePSLMC locking troubles: MC/FSS servo unstable

The MC seemed to be losing lock recently quite a bit. I noticed that the PC Drive RMS signal was red.

This means that the high frequency drive to the Pockels cell was too high by a factor of 2-3 and sometimes saturating and breaking the lock.

I fiddled with the gains on the FSS screen until this value went down. It looks like there is some kind of high Q oscillation; it takes a couple minutes for the PC Drive RMS to settle to its new position after changing the gains.

The attached trend plot show the last 2 hours. The mean is now back to ~1 V and seems OK. We should really examine the FSS or MC error point spectra with the RF analyzer while exploring this gain space.

Attachment 1: Untitled.png
Untitled.png
  8423   Mon Apr 8 16:37:26 2013 SteveUpdate40m UpgradingETMY enclosure wall transmission

 

 Enclosure cover #1 transmission measured  in 1064 nm,  156 mW,  P polarization and beam size ~ 1 mm

As condition:  fully assembled, protective layer removed, tinted- adhesive activated on yellow acrylic on top of each other.

T = 1.2 % in 20 minutes exposure test. This agrees with the test measurement of 6-18-2012

There is a reflected  2-3 cm circular glare that is barely visible on sensor card. It is well below 1 mW level

As we are installing the NPRO with ~350 mW of power we have to address what additional shield should be installed.

The June 2012 test with 1W power burned through of the 3 layer  IR coated films in 3-4 hours.

We 'll use Aluminum shields in the high power path till we come up with better solution.

  8422   Mon Apr 8 10:19:46 2013 JenneUpdatePSLLSC left enabled

 

 Note: The TRY PD isn't installed and normalized properly yet, so the IFO OVERVIEW screen indicates lock for the Yarm constantly, which is not true.  Hopefully in the next day or so the screen will be back to telling the truth.

Also, the LSC Locking was left ENABLED (presumably over the weekend).   This is not so good.  It can kick optics around, so we should all take a look when we walk through the control room, and if no one is locking, please disable the LSC master switch. 

  8421   Mon Apr 8 08:05:41 2013 SteveUpdatePSLPMC locked

The PMC locked manually. MC  grabbed lock instantaniously

  8420   Sun Apr 7 20:49:19 2013 ZachUpdateGeneralRestarted elog

with the script, as it was down.

  8419   Sat Apr 6 09:21:36 2013 ranaUpdate40m UpgradingNot a fan of the new plastic box yet

 

 1) We still need to drill and install the thumbscrew latches which secure the lids to the table. We cannot use the tables as an acoustic enclosure with loose lids.

2) For the camera issue, the idea is to put the longpass filters on the front of the cameras: then they are only sensitive to light with wavelength > 800 nm.

3) Whenever any interferometer work is happening the light switches must be in the positions which have been marked on them (and which most everyone ignores foolishly). We have never been insensitive to the room lights; the black table enclosures just give us a false sense of security. Room lights impact the interferometer noise.

  8418   Fri Apr 5 01:28:56 2013 ManasaUpdate40m UpgradingEndtable upgrade for auxiliary green laser : populating the table

I started populating the end table; the TRY path to start with. I found that I need to redo the cables/electronics layout around the table as we have only one cable feedthrough hole with the new box right now. I need another hand with this and will have Annalisa help me tomorrow.

P.S. I misaligned PRM and restored ETMY to get TRY flashes. I tweaked ETMY to see strong TEM00 flashes.

Old slider positions on medm screen in case we need to restore them:

        TT1    TT2     ITMY    ETMY

P   -1.3586  0.8443   0.9114   -3.7693

Y   0.3049   1.1507  -0.2823  -0.2761

 

  8417   Fri Apr 5 01:18:34 2013 ManasaUpdate40m UpgradingNot a fan of the new plastic box yet

Quote:

Quote:

 Enclosure is at the east end. It has it's bottom o-ring in place. It will be ready for optics tomorrow around 5pm

I have to shim out the enclosure, finish leveling the table and cut surgical tubing O-ring for the top.

 

 Glued surgical latex tubing with super glue into O-ring shape. The existing in place tubing K-100, OD 0.125" (actual size 0.140"), wall 0.031", ID 0.062".

I have just found out that tolerances on tubing OD are + - 0.026" by the manufacturer. I'm getting larger tubing for better fit.

The table is ready for optics.

Things left to do:

1, finalize o-ring size  2, finish cable feedthrough  3, finalize window connection 4, IR-Thermashield strips for bridge sides

 

While I did think that the plastic boxes were cool; I am not happy with the new enclosure on several aspects as I am setting up the TRY PDs/camera.

1. I strongly feel the o-rings should be glued/fixed to the plastic box atleast at some points. Even if we replace the current thin one with a thicker tubing, it takes forever to get the o-ring into the groove when it slips out. Also if it slips from the groove, it falls through the optical path and there are chances of burning the tubing when the NPRO is ON.

2. With the transparent tables, the cameras are sensitive to the room lights. I had to switch off the room lights/use ND filters to see a nice beam at the TRY camera.

3. The lids are heavy...might be a good way to train....but removing and putting them back will definitely increase the pain in getting the green aligned to the arms atleast until we have the PZTs set up.

  8416   Thu Apr 4 17:23:05 2013 SteveUpdate40m Upgrading ETMY optical table & enclosure is ready for optics

Quote:

 Enclosure is at the east end. It has it's bottom o-ring in place. It will be ready for optics tomorrow around 5pm

I have to shim out the enclosure, finish leveling the table and cut surgical tubing O-ring for the top.

 

 Glued surgical latex tubing with super glue into O-ring shape. The existing in place tubing K-100, OD 0.125" (actual size 0.140"), wall 0.031", ID 0.062".

I have just found out that tolerances on tubing OD are + - 0.026" by the manufacturer. I'm getting larger tubing for better fit.

The table is ready for optics.

Things left to do:

1, finalize o-ring size  2, finish cable feedthrough  3, finalize window connection 4, IR-Thermashield strips for bridge sides 5, replace bridge support post with solid one

 

  8415   Thu Apr 4 14:37:15 2013 JamieConfigurationElectronicsputting together a 110 MHz LSC demod board

I'm having Steve order the following:

2x  SXBP-100+
2x  SCLF-135+
2x  PSCQ-2-120+

If you want him to add anything to the order let him know ASAP.

  8414   Thu Apr 4 13:39:12 2013 Max HortonUpdateSummary PagesGraph Limits

Graph Limits: The limits on graphs have been problematic.  They often reflect too large of a range of values, usually because of dropouts in data collection.  Thus, they do not provide useful information because the important information is washed out by the large limits on the graph.  For example, the graph below shows data over an unnecessarily large range, because of the dropout in the 300-1000Hz pressure values.

Time series data from frames

The limits on the graphs can be modified using the config file found in /40m-summary/share/c1_summary_page.ini.  At the entry for the appropriate graph, change the amplitude-lim=y1,y2 line by setting y1 to the desired lower limit and y2 to the desired upper limit.  For example, I changed the amplitude limits on the above graph to amplitude-lim=.001,1, and achieved the following graph.

Time series data from frames

The limits could be tightened further to improve clarity - this is easily done by modifying the config file.  I modified the config file for all the 2D plots to improve the bounds.  However, on some plots, I wasn't sure what bounds were appropriate or what range of values we were interested in, so I will have to ask someone to find out.

Next:  I now want to fix all the funny little problems with the site, such as scroll bars appearing where they should not appear, and graphs only plotting until 6PM.  In order to do this most effectively, I need to restructure the code and factor it into several files.  Otherwise, the code will not only be much harder to edit, but will become more and more confusing as I add on to it, compounding the problems that we currently have (i.e. that this code isn't very well documented and nobody knows how it works).  We need lots of specific documentation on what exactly is happening before too many changes are made.  Take the config files, for example.  Someone put a lot of work into them, but we need a README specifying which options are supported for which types of graphs, etc.  So we are slowed down because I have to figure out what is going on before I make small changes.

To fix this, I will divide the code into three main sectors.  The division of labor will be:
- Sector 1: Figure out what the user wants (i.e. read config files, create a ConfigParser, etc...)
- Sector 2: Process the data and generate the plots based on what the user wants
- Sector 3: Generate the HTML

  8413   Thu Apr 4 10:46:54 2013 KojiSummaryLSCREFL55 error signals

Beautiful double peaks. I don't see the triple zero-crossings. Is this because you adjusted the phase correctly (as predicted)?

Don't you want to have a positive number for POP22? Should we set the demod phase in the configuration script for the positive POP22, shouldn't we?

  8412   Thu Apr 4 10:32:42 2013 GabrieleSummaryLSCREFL55 error signals

The attached plot shows that also the behaviour of the REFL11 and 55 signals is qualitatively equal to the simulation outcome.

Attachment 1: prmi_refl_signals.pdf
prmi_refl_signals.pdf
  8411   Thu Apr 4 10:12:55 2013 GabrieleSummaryLSCPOP22 and POP110

I had a look at the POP110 signal, with the PRMI flashing.

1) The LSCoffset script does not zero any more POP22_I_ERR offset. I did it by hand

2) The gain of POP22 is changed a lot, as well as the sign: now sidebands are resonant when POP22_I is negative

3) POP110 seems to deliver good signals. The plot attached shows that when we cross the sideband resonance, there is a clear splitting of the peak. If we rely on the simulations I posted in entry 8401, the full width at half height of the POP_22 peak is of the order of 5 nm. Using this as a calibration, we find a splitting of the order of 7 nm, which is not far from the simulated one (5 nm)

Attachment 1: pop110_2.pdf
pop110_2.pdf
  8410   Wed Apr 3 23:22:20 2013 ranaUpdateoptical tablesOptical Table Toolboxes Update

 There are some tips for how to appy nail polish on YouTube from MKNails and MissJenFABULOUS. Their tips on how to prepare the site for a strong bonding strength are probably helpful for our gold/nickel coated tools. For chrome tools we may need to abrade the surface with a stone or fine sandpaper for it to take the layer better. IF the YouTube videos don't do it for you, then I suggest contacting Tom Evans at LLO to find out what kind of nail polish he uses.

  8409   Wed Apr 3 22:26:51 2013 ranaConfigurationElectronicsputting together a 110 MHz LSC demod board

 For the 110 MHz demod boards, we would ideally have a plugin bandpass filter. If you have some specs in mind, you can email mini-circuits or pulsar microwave about making a custom part; its not too expensive usually.

For the meantime, you should remove the onboard one and replace with a combination of low/high pass filters from Mini-Circuits. If you put a SLP-150 and a SHP-100 in series, the insertion loss should be less than 1 dB.

I think the ERA amps are OK for now, but they die with time, so they just need to be tested and replaced if necessary.

  8408   Wed Apr 3 19:01:06 2013 AlbertUpdateoptical tablesOptical Table Toolboxes Update

Quote:

Quote:

A heavy duty plastic box is the likeliest candidate for the optical table toolbox. It measures 5 9/16 in. x 11 5/8 in. x 4 5/8 in. and fits all the tools comfortably. ( http://www.mcmaster.com/#plastic-bin-boxes/=m4yh4m  ,  under Heavy Duty Plastic Bin Boxes)

The list of tools has been updated to include a pen and a wire cutter as well as everything previously stated.

In addition, Steve has recommended that boxes should be secured to the walls or surfaces near the optical tables as opposed to the optical tables themselves, as to keep the tables from wobbling when tools are being exchanged.

A diagram of tentative box placements will go out soon.

 No, the small boxes must be attached to the optical tables. They won't be heavy enough to change the table tilt.

Also, all tools must be color coded according to the optical table using the 3M Vinyl table color code:

http://www.3m.com/product/images/Vinyl-Electrical-Color-Tape-300.jpg

 Ok.

So the new tentative plan on the boxes is to bolt them (magnetic strips were proposed but overruled on the grounds that they're not strong enough to withstand being knocked down by accidents).

The boxes are going to be a mix of the Thorlabs Benchtop Organizer (http://www.thorlabs.com/thorProduct.cfm?partNumber=BT17) and the original box. The box will have a region covered in mesh, so tools can be placed and held there. The box will also have a spacer at the bottom, with another mesh right above it, lined up. However, this double-mesh will only cover half of the box. The other half of the box will be compartmentalized to hold things such as screws, connectors, etc. I will talk to Steve about building the boxes.

Also, using nail-polish to coat the Allen wrenches is not going to work. Nail polish does not stick easily enough. The tentative new plan is oil paint, but this is to be reviewed.

Finally, the diagram with the placement of the boxes relative to the optical tables has been put on paper, but needs to be computerized so it's easier to read. This will be done as soon as possible.

  8407   Wed Apr 3 18:41:22 2013 JamieConfigurationElectronicsputting together a 110 MHz LSC demod board

This SCPQ-150+, which is surface mount, might also work in place of the PSCQ-2-120, which is through-mount.  Would need to be reconciled with the board layout.

  8406   Wed Apr 3 18:27:03 2013 KojiUpdate QPD Voltage Regulators

Breadboards may not be suitable for a reliable work. Why don't you switch to any protoboard and real soldering?

  8405   Wed Apr 3 18:22:00 2013 JenneUpdateElectronicsPOP110 re-implemented

I have re-implemented POP110.  The cable coming from the AS110 diode is disconnected, labeled, and sitting in the cable tray next to the LSC rack. 

Now the POP diode path is:

 

Thorlabs 10CF ----many meters of heliax cable-----> Bias Tee ------> RF amplifier ------> Splitter ------> Bandpass 21.7MHz --------> POP22 demod board

                                                                                                                   |                                                                                    |

                                                                                                                   |                                                                                    |

                                                                                                                   V                                                                                  V

                                                                                                            POP DC                                                                        High pass 100MHz

                                                                                                                                                                                                        |

                                                                                                                                                                                                        |

                                                                                                                                                                                                       V

                                                                                                                                                                                              Lowpass 150MHz

                                                                                                                                                                                                       |

                                                                                                                                                                                                       |

                                                                                                                                                                                                      V

                                                                                                                                                                                        POP110 demod board

  8404   Wed Apr 3 17:40:18 2013 JamieConfigurationElectronicsputting together a 110 MHz LSC demod board

I started to look into putting together a 110 MHz demod board to be used as POP110 (see #8399).

We have five spare old-skool EuroCard demod boards (LIGO-D990511).  From what I gather (see #4538, #4708) there are two modifications we do to these boards to make them ready for prime time:

  • appropriate LP filter at PD RF input (U5 -> MC SCLF-*)
  • swap out T1 transformer network with a commercial phase shifting power splitter (MC PQW/PSCQ)

#4538 also describes some other modifications but I'm not sure if those were actually implemented or not:

  • removal of the attenuator/DC block/ERA-5 amp sections at the I/Q outputs
  • swap ERA-5 amp with "Cougar"(?) amp at LO input.

What we'll need for a 110 demod:

I'll scrounge or order.

  8403   Wed Apr 3 16:03:59 2013 ChloeUpdate QPD Voltage Regulators

The voltage regulator on the QPD breadboard seems to be having problems... yesterday Eric helped me debug my circuit and discovered that the +12V regulator was overheating, so we replaced it. Today, I found that the -12V regulator was also doing the same thing, so I replaced it. However, it's still overheating. We checked all of the setup for the power regulators yesterday, so I'm not sure what's wrong.

I've also noticed that not all the connections on the breadboard that I've been using seem to work - I may search for a new breadboard in this case. Need to check I'm not doing something stupid with that.

  8402   Wed Apr 3 15:00:24 2013 JamieSummaryElectronicsSorensen supplies in LSC rack (1Y2)

I investigated the situation of the two Sorensen supplies in the LSC rack (1Y2).  They are there solely to supply power to the LSC LO RF distribution box.  One is +18 V and the other is +28 V.  All we need to do is make a new longer cable with the appropriate plug on one end (see below), long enough to go from the bottom of the 1Y3 rack to the top of 1Y2, and we could move them over quickly.  Some sort of non-standard circular socket connector is used on the distribution box:

20130403_141842.jpg

It could probably use thicker conduction wire as well.

If someone else makes the cable I'll move everything over.

  8401   Wed Apr 3 14:46:17 2013 GabrieleSummaryLSCError signal simulation in PRMI

Here is a summary of a simulation of the error signal behavior in the PRMI configuration. The main parameters are:
L_PRC = 6.7538 m
Schnup = 0.0342 m
fmod1 = 11.065399e6 Hz
fmod2 = 5 * fmod1

prmi_prclsweep_pop22.pngprmi_prclsweep_pop110.png

These two plots shows the response of the POP22 and POP110 signals (in almost arbitrary units) to a PRCL sweep around the resonance. The splitting of the 55 sideband peaks is well visible in the second plot. It is due to the fact that the 55MHz sidebands are not perfectly matched to the PRC length

prmi_michsweep_pop22.pngprmi_michsweep_pop110.png

The same thing when sweeping MICH. The peaks are wider and it is not possible to see the splitting.

prmi_prclsweep_errsig_not_tunedphase.pngprmi_michsweep_errsig_not_tunedphase.png

These are the error signals (REFL11_I/Q and REFL_55_I/Q) as a function of the PRCL (left) and MICH (right) sweep. Here the demodulation phases are not properly tuned. This is just to show that when the phase is wrong, you can get multiple zero crossings (in this case only in the Q signals, but in general also in I) close to resonance.

prmi_prclsweep_errsig_tunedphase.pngprmi_michsweep_errsig_tunedphase.png

If the phases are tuned in order to maximize the slope of the I signals with respect to PRCL, one gets these "optimized phase" responses. It is that the phase does not correspond to the one that makes the PRCL peak to peak signal small in Q. The Q signals are indeed flat around resonance for a PRCL motion, but they deviate quite a lot from zero when moving more far from resonance. Moreover, both the REFL_55 error signals (I for PRCL and Q for MICH) are crossing again zero at two additional positions, but those are quite far from the resonance point.

prmi_prclsweep_triggering.pngprmi_michsweep_triggering.png

These plots just show the PRCL and MICH error signals together with the POP22 and POP110 signals, to give an idea of the level of triggering that might be needed to be inside the linear range. It seems that if we trigger on POP22 when using the REFL55 signal we loose a bit of linear range, but not that much.

prmi_prclsweep_linearized_signals.pngprmi_michsweep_linearized_signals.png

If you reached this point it means you're really interested in this topic, or maybe you have nothing better to do... However, this plot shows the effect of linearization of the error signal, obtained dividing them by the proper POP22/110 signal. The linear range is increased, but unfortunately for the 55 signals, the additional zero crossing I was mentioning before creates two sharp features. Those are however quite outside the triggering region, so they should not be harmful.

Attachment 1: prmi_michsweep_pop22.png
prmi_michsweep_pop22.png
Attachment 2: prmi_michsweep_pop110.png
prmi_michsweep_pop110.png
  8400   Wed Apr 3 14:45:34 2013 JamieUpdateComputersupdated EPICS database (channels selected for saving)

Quote:

I modified /opt/rtcds/caltech/c1/chans/daq/C0EDCU.ini to include the C1:LSC-DegreeOfFreedom_TRIG_MON channels.  These are the same channel that cause the LSC screen trigger indicators to light up. 

I vaguely followed Koji's directions in elog 5991, although I didn't add new grecords, since these channels are already included in the .db file as a result of EpicsOut blocks in the simulink model.  So really, I only did Step 2.  I still need to restart the framebuilder, but locking (attempt at locking) is happening.

The idea here is that we should be able to search through this channel, and when we get a trigger, we can go back and plot useful signals (PDs, error signals, cotrol signals,....), and try to figure out why we're losing lock. 

Rana tells me that this is similar to an old LockAcq script that would run DTT and get data.

EDIT:  I restarted the daqd on the fb, and I now see the channel in dataviewer, but I can only get live data, no past data, even though it says that it is (16,float).  Here's what Dataviewer is telling me:

Connecting to NDS Server fb (TCP port 8088)
Connecting.... done
read(); errno=0
LONG: DataRead = -1
No data found

read(); errno=9
read(); errno=9
T0=13-03-29-08-59-43; Length=432010 (s)
No data output.
 

I seem to be able to retrieve these channels ok from the past:

controls@pianosa:/opt/rtcds/caltech/c1/scripts 0$ tconvert 1049050000
Apr 03 2013 18:46:24 UTC
controls@pianosa:/opt/rtcds/caltech/c1/scripts 0$ ./general/getdata -s 1049050000 -d 10 --noplot C1:LSC-PRCL_TRIG_MON
Connecting to server fb:8088 ...
nds_logging_init: Entrynds_logging_init: Exit
fetching... 1049050000.0
Hit any key to exit: 
controls@pianosa:/opt/rtcds/caltech/c1/scripts 0$ 

Maybe DTT just needed to be reloaded/restarted?

  8399   Wed Apr 3 03:18:23 2013 JenneUpdateLockingPRMI locked with REFL55 I&Q for more than 1 minute

[Rana, Gabriele, Jenne]

We have now locked the PRMI using REFL55 I&Q for more than one minute!!!!!

PRMI_sb_80sec_2Apr2013.png

This isn't really the most useful plot as is, but it was created using:

/opt/rtcds/caltech/c1/scripts/general/getdata C1:LSC-POP22_I_ERR_DQ C1:LSC-REFL55_I_ERR_DQ C1:LSC-REFL55_Q_ERR_DQ C1:LSC-MICH_IN1_DQ C1:LSC-MICH_OUT_DQ C1:LSC-PRCL_IN1_DQ C1:LSC-PRCL_OUT_DQ -d 80 -s 1049013520 -c

This is just one of several long lock stretches.  If I can get the TRIG_MON channels to be saved, we can automatically (versus my by-hand search) find lock stretches and make this kind of plot.  Although we want them saved in some raw format so we can zoom in on selected axes, I think.  This might require some python-fu from Jamie, or learning of python-fu for Jenne.

The secret sauce:

* The big key was changing REFL55's phase.  It was -4 when we looked at the I&Q signals, and minimized the PRCL information in the Q-phase.  We were able to get short lock stretches with this.  During these stretches, Rana changed the REFL55 phase until the lock sounded (audibly) quieter.  The final phase we settled on was +26. As we changed the phase, the lock stretches got longer and longer. 

* We also tweaked up the POP22 phase.  It was close from our previous efforts of looking at non-locked time series, but we perfected it by minimizing the signal in the Q-phase during lock stretches.  We also found that it drifted (according to this method) by ~5 degrees over ~half an hour (I don't remember the exact time between our phase tunings).

* POP22's low pass filters (both options, ELP10 and ELP50) must be OFF for any lock to be acquired.  Turning on either filter prevents locking.

* Normalization helped a lot.  Without normalization we weren't really able to catch any locks, certainly not of any significant length. (0.004, using POP22I, for both MICH and PRCL). 

* Parameters:

       ** Normalization:  use POP22I for both MICH and PRCL, value = 0.004

       ** Input matrix:  MICH with REFL55Q, value = 0.01; PRCL with REFL55I, value = 0.01  (we used the small number in the matrix so our servo gains weren't too tiny).

       ** POP22 lowpass filters OFF

       ** Analog whitening OFF for REFL55, POP22.

       ** Analog gain for REFL55 I&Q = 27 dB

       ** Analog gain for POP22 I&Q = 15 dB

       ** Output matrix:  MICH with -1 to ITMX, +1 to ITMY.  PRCL with +1 to PRM.

       ** Servo gains:  PRCL = 0.75;  MICH anywhere between -3 and -20.  Best in the -8 to -15 range.

       ** Vio2 filters in ITMX, ITMY, PRM (all actuated-on mirrors) were OFF. (Still need to lower the Q on these so they don't ring).

      ** PRCL and MICH triggering on POP22I.  The trigger-off was always 20, but the trigger-on changed throughout the night from ~170 to ~50.  I think 130 was a trigger value for at least some of the long-time locks.

      ** Low frequency seismic was small (i.e. no anomalous 0.1 Hz - 1 Hz noise) during successful lock times.  (Not to say it must be low, but it was low when we were able to lock for long stretches).

Things we had looked at and thought about throughout the evening:

* Oplev calibration.  See elog 8391 and 8393.  Optimized BS and PRM to reduce yaw angular motion.

* Actuators all functioning as expected.  We checked transfer functions of MICH_OUT/MICH_IN1 for locking with different optics, to ensure that at high frequency the response was 1/f^2.  Also, we locked MICH with (a) both ITMs, (b) BS, (c) ITMX and (d) ITMY.  We locked the PR-ITMY half-cav with (a) PRM and (b) ITMY.  We locked the PR-ITMX half-cav with (a) PRM and (b) ITMX.  Thus, we conclude that all of the PRMI-related optics are functioning as expected.

* Realigned REFL55 beam onto PD.  It was clipping a bit, so the DC power wasn't steady (when ITMs were misaligned, PRM aligned).  After alignment, the DC power as seen on a 'scope was much smoother.

* Turning off the limiters for the MICH and PRCL control signals allowed us to hear a high-pitched whine.  From looking at the time series, it's predominantly in MICH_OUT.  Rana speculates that perhaps the normalization is causing the UGF to wander temporarily to an unstable place.  For a time there was a high-Q peak between 500 and 600Hz, but reducing the gain (of MICH?) eliminated that.  Then we heard several times, irrespective of gain setting, the ~400Hz broad peak (I say broad because I was able to see it on DTT looking at the error and control signals, and it spanned +/-100Hz).

Things to investigate:

* Is there a good reason that we should switch to triggering on POP110, rather than the current POP22?  From Gabriele, Jamie and my Finesee/Mist modelling last week, without the arms, the 11MHz and 55MHz resonate at different PRC lengths.  If this difference is very small, then we are fine, but if the difference is large, it could be causing trouble - we're trying to catch the lock at the linear part of the 55MHz signal, but if that does not coincide with the linear part of the 11MHz signal, we're doing the wrong thing.  

* For the POP normalization, should we be using the amplitude or the power ( POP22 or sqrt(POP22) )?  Why?  Look at this with a modelling sweep and/or analytically.

* Look at different noise sources, potentially sensing noise, coil actuator noise,.....  We should check these out, and make sure we're not limited by anything obvious. 

Other To-Do:

* Make a "restore" medm screen, rather than restore script.  IFO Configure restore script can pull in values from the screen (EPICS values).  One screen per configuration. 

* Get TRIG_MON signals saved, write script to search for triggered lock times (between given gps times), then plot interesting signals for just before lock, during lock, and until just after a lockloss.

  8398   Wed Apr 3 01:32:04 2013 JenneUpdateComputersupdated EPICS database (channels selected for saving)

I modified /opt/rtcds/caltech/c1/chans/daq/C0EDCU.ini to include the C1:LSC-DegreeOfFreedom_TRIG_MON channels.  These are the same channel that cause the LSC screen trigger indicators to light up. 

I vaguely followed Koji's directions in elog 5991, although I didn't add new grecords, since these channels are already included in the .db file as a result of EpicsOut blocks in the simulink model.  So really, I only did Step 2.  I still need to restart the framebuilder, but locking (attempt at locking) is happening.

The idea here is that we should be able to search through this channel, and when we get a trigger, we can go back and plot useful signals (PDs, error signals, cotrol signals,....), and try to figure out why we're losing lock. 

Rana tells me that this is similar to an old LockAcq script that would run DTT and get data.

EDIT:  I restarted the daqd on the fb, and I now see the channel in dataviewer, but I can only get live data, no past data, even though it says that it is (16,float).  Here's what Dataviewer is telling me:

Connecting to NDS Server fb (TCP port 8088)
Connecting.... done
read(); errno=0
LONG: DataRead = -1
No data found

read(); errno=9
read(); errno=9
T0=13-03-29-08-59-43; Length=432010 (s)
No data output.

 

  8397   Tue Apr 2 23:14:02 2013 ranaUpdateLSC22/110MHz path for POP

Quote:

We could not find a power supply slot for the amplifiers on the LSC rack. We had to put a temporary power supply in contradiction to our 'no temporary power supply' policy.

 After 1 month, its hard to imagine that this could not have been fixed by putting in a proper fuse and fuse block. I will remove this tomorrow if I still find it this way in the bottom of the rack.

There are also 2 Sorensen switching supplies in the bottom of the LSC rack (with all of our sensitive demod boards). These should also be moved over to the old 'digital' LSC rack tomorrow for the post meeting lab cleanup.

Use fuse blocks with fuses with appropriate ampacity.

  8396   Tue Apr 2 22:39:17 2013 AnnalisaUpdateAuxiliary lockingBeat note between "Alberto" NPRO laser and PSL laser

 

 The beat note between the PSL laser and the "Alberto" NPRO laser has been measured. In particular, for each PSL temperature, more than one Aux laser frequency has been found.

The second of the three curves seems to be more stable than the other two, even if a "step" trend can be found in all of them (maybe due to the frequency change of the NPRO laser as a function of the crystal temperature continuous tuning, as mentioned in the previous elog). This is the reason why the points are not perfectly aligned, and the errors on the fit parameters are so big.

 

 

Attachment 1: Beat_note_3col.jpg
Beat_note_3col.jpg
ELOG V3.1.3-